what is two more than a number is the same as 16 dereased by 6 times the number
Answer by checkley77(12569)

X+2=16-6X
X+6X=16-2
7X=14
X=14/7
X=2 ANS.
PROOF:
2+2=16-6*2
4=16-12
4=4